# Context-Aware Replanning with Pre-explored Semantic Map for Object Navigation

### v1

@@ Acknowledgments
To further demonstrate the real-time applicability of our method, we measured the time requirements in the OpenMask3d experiment setting. We assumed that object features are precomputed and stored in the maps, with uncertainty measures such as entropy and KL divergence computed on the fly. Figures 5 and 5 present the results under different top-k strategies. For all uncertainty measures, both top-k selection criteria and all tested values of $k$, a retrieval attempt typically takes less than 15 milliseconds. Considering that real-world physical navigation involves seconds or minutes of path-following interactions after determining a destination, we believe that the millisecond-level computational overhead introduced by our method is negligible.
@@
Figure 4: Retrieval Latency for Top-k Confidence.
Figure 5: Retrieval Latency for Top-k Category.

### v2

@@ Acknowledgments
Main Augmentation:
To enable multi-view uncertainty calculation, we augmented OpenMask3D not to take the average of the multi-view features, but to record features from all visible views for each object mask.
Let:
  • $P$ be the scene point cloud.
  • $I=\{I_{1},I_{2},\ldots,I_{n}\}$ be the set of posed RGB images.
  • $V_{i}$ be the set of views where object $m_{i}$ is visible.
$F_{ij}=\text{CLIP}(\text{crop}(I_{j},m_{i})),\quad\forall I_{j}\in V_{i}$
where $F_{ij}$ is the feature vector for object $m_{i}$ from view $I_{j}$.
$F_{i}=\{F_{ij}\mid I_{j}\in V_{i}\}$
@@
Channel-average Feature Standard Error for VLMaps:
Recall that for each candidate $c_{i}$ with multiple views, we compute the standard error $SE(c_{i})$:
$SE(c_{i})=\frac{1}{d}\sum_{k=1}^{d}\frac{\sigma_{k}}{\sqrt{m_{i}}}$
where $\sigma_{k}$ is the standard deviation of the $k$-th feature across all views, $d$ is the dimensionality of the feature vector, and $m_{i}$ is the number of views for candidate $c_{i}$.
And in the case of VLMaps, we incorporate the distance weighting [1, 5] information of each feature which is utilized when building the map, that is we replace the standard deviation and the sample size with the weighted version:
$SE^{w}(c_{i})=\frac{1}{d}\sum_{k=1}^{d}\frac{\sigma^{w}_{k}}{\sqrt{m^{eff}_{i}}}$
$m^{eff}_{i}=\frac{(\sum_{i=1}^{n}w_{i})^{2}}{\sum_{i=1}^{n}{w_{i}^{2}}}$
Finally, we choose the candidate $c^{*}$with the lowest standard error.
@@
Figure 5: Latency for Top-k Confidence.
Figure 6: Latency for Top-k Category.
To further demonstrate the real-time applicability of our method, we measured the time requirements in the OpenMask3d experiment setting. We assumed that object features are precomputed and stored in the maps, with uncertainty measures such as entropy and KL divergence computed on the fly. Figures 6 and 6 present the results under different top-k strategies. For all uncertainty measures, both top-k selection criteria and all tested values of $k$, a retrieval attempt typically takes less than 15 milliseconds. Considering that real-world physical navigation involves seconds or minutes of path-following interactions after determining a destination, we believe that the millisecond-level computational overhead introduced by our method is negligible.